Aqueous lead (II) sulfate combines with aqueous magnesium bromide to form what?

Silver iodide is mixed with potassium carbonate to form what?

Aqueous ammonium phosphate reacts with aqueous mercury (II) sulfate to form what?
1
Expert's answer
2020-03-30T10:50:55-0400

MgBr2+PbSO4 = MgSO4+Br2Pb

4AgI + 2K2CO3 = 4Ag + 4 KI + 2CO2+ O2

2(NH4)3PO4 + 3 HgSO4 = Hg3(PO4)2 + 3(NH4)2SO4